Variable Costing Unit Product Cost and Income Statement

Chuck Wagon Grills, Inc., makes a single product - a handmade specialty barbecue grill that it sells for $210. Data for last year's operations follow:

Units in beginning inventory 0
units produced 20,000
units sold 19,000
units in ending inventory 1,000

Variable costs per unit:
Direct materials $50
Direct Labor 80
Variable manufacturing overhead 20
Variable selling and admin. 10

Total variable cost per unit $160

Fixed Costs:
Fixed Manufacturing Overhead $700,000
Fixed Selling and Admin. 285,000
Total Fixed Costs $985,000

1. Assume the company uses variable costing. Compute the unit product cost for one barbecue grill.
2. Assume the company uses variable costing. Prepare a contribution format income statement for the year.
3. What is the company's break-even point in terms of of the number of barbecue grills sold?
Answer #1
1)Under varible costing menthod:
units sold = 19000 units
total fixed cost = $985000
per unit = 985000 /19000 =$51.84
Varible cost = $160 per unit
Total product cost per unit = 51.84+160=$211.84
2) under vairable costing menthod :
Contribution per unit  = $210-$160 = $50
Income stataemetn:
sales 3990000
less: Variablecost 3040000
Contribution 950000
less: fixed cost 985000
loss -35000
3)
Break even point = Fixed cost / contribution per unit
= $985000 / $50
= 19700 units
Under Absobtion cost :
Product cst :
here unit cost is calicuated :
Toal varible cost + fixecd cost
$160 +[ 985000 /21000]
+$160+40.9
= $200.9
2) income statement :
sale 3990000
add:ending inventory 211840
lee: cost of good sold 4018000
gross loss -172900
here unit cost is calicuated :
Toal varible cost + fixecd cost
$160 +[ 985000 /21000]
+$160+40.9
= $200.9
Answer #2

uploaded image

uploaded image

Break-even point in sales dollars = [$985,000 / ($210 - $160) / $210]

Break-even point in sales dollars = $985,000 / 0.238095 = $4,137,000
Answer #3
Concepts and reason

Variable Costing: Variable costing is a standard whereby variable expenses are collected to costs units and the fixed expenses traceable to the related period is written off in full beside the contribution for that period.

Contribution Margin Income Statement: Contribution margin income statement highlights the cost behavior as variable costs and fixed costs and shows the contribution margin of the company.

Contribution Margin Ratio: The difference between the sales revenue and the variable costs is called contribution margin. Contribution margin ratio calculated by dividing the contribution margin by selling price. Revenue required to earn the target net income can be determined by using this contribution margin ratio.

Break Even Point in units: Breakeven point in units (BEP) is the point of sales at which there is no profit or loss. It means the total units equal the total revenue generate at this point.

Fundamentals

Selling price: Selling price is the price at which each product is sold in the market.

Variable cost: These costs vary with the number of units produced or for the services provided. For example, the labor costs increase if the number of labor hours is increased, and the labor costs decrease if the number of labor hours is decreased.

Fixed cost: These are the costs which remain constant throughout the process of manufacturing or for the services rendered. They are incurred irrespective of number of units produced. For example, factory rent. The rent should be paid, if the factory produces 100 units or 1,000 units or if no units are produced.

Formula for Variable cost of Goods sold as follows:

VariablecostofGoodssold={Openninginventory+UnitsproducedEndinginventory}{\rm{Variable cost of Goods sold = }}\,\left\{ \begin{array}{l}\\{\rm{Openning inventory + Units produced}}\\\\{\rm{ }} - \,{\rm{Ending inventory}}\\\end{array} \right\}

Formula for Break –Even point in units as follows:

BreakEvenpoint(numbers)}=Fixedcosts(SellingpriceperunitVariablecostperunit)\left. \begin{array}{l}\\{\rm{Break Even point}}\\\\{\rm{ (numbers)}}\\\end{array} \right\}{\rm{ = }}\,\frac{{{\rm{Fixed costs}}}}{{\left( \begin{array}{l}\\{\rm{Selling price per unit }} - \,\\\\{\rm{Variable cost per unit}}\\\end{array} \right)}}

(1)

Compute unit product cost for one barbecue grill sold under variable costing as follows:

Particulars
Direct materials
Direct labor
Variable manufacturing overhead
Unit product cost for one barbecue grill
Per unit
$

Hence, Unit product cost for one barbecue grill is $150.

(2)

Working note:

Calculate variable cost of goods sold units as follows:

VariablecostofGoodssold={Openninginventory+UnitsproducedEndinginventory}=0+20,0001,000=19,000Units\begin{array}{c}\\{\rm{Variable cost of Goods sold = }}\,\left\{ \begin{array}{l}\\{\rm{Openning inventory + Units produced}}\\\\{\rm{ }} - \,{\rm{Ending inventory}}\\\end{array} \right\}{\rm{ }}\\\\{\rm{ = 0}}\,{\rm{ + }}\,{\rm{20,000}} - \,1,000\\\\ = 19,000\,{\rm{Units}}\\\end{array}

Calculate contribution format income statement for CWG incorporation as follows:

Sales
CWG Incorporation
Contribution Format Income Statement
Particulars
Units (A) Per Unit (B)
19,000 $210
Less: Variable co

Hence, Net Operating loss is $35,000.

Calculate Break – Even Point (numbers).

BreakEvenpoint(numbers)}=Fixedcosts(SellingpriceperunitVariablecostperunit)=$985,000($210$160)=$985,000$50=19,700Units\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Break Even point}}\\\\{\rm{ (numbers)}}\\\end{array} \right\}{\rm{ = }}\,\frac{{{\rm{Fixed costs}}}}{{\left( \begin{array}{l}\\{\rm{Selling price per unit }} - \,\\\\{\rm{Variable cost per unit}}\\\end{array} \right)}}\\\\ = \,\,\frac{{\$ 985,000}}{{\left( {{\rm{\$ 210 }} - \,\$ 160\,\,} \right)}}\\\\ = \,\frac{{\$ 985,000}}{{\$ 50}}\\\\ = 19,700\,{\rm{Units}}\\\end{array}

Hence, Break-even point (numbers) is 19,700.

Ans: Part 1

Unit product cost for one barbecue grill is $150.

Part 2

Sales
CWG Incorporation
Contribution Format Income Statement
Particulars
Units (A) Per Unit (B)
19,000 $210
Less: Variable co

Part 3

Breakeven point (numbers) of barbecue grills sold is 19,700.
Answer #4

(1)

Compute the unit product cost for one barbecue grill:

Picture 1

Therefore, the unit production cost is

(2)

Compute the contribution income statement for the year:

Picture 1

Therefore, the net operation loss for the year is

(3)

Compute break-even point in terms of number of barbecue grills sold:

Therefore, the breakeven points in units are.

The company’s break-even point in sales dollars is. This is computed by dividing total fixed cost by the contribution margin.